OpenStudy (anonymous):

quadratic formula can be used to solve the fourth degree equation ax^4+bx^2+c=0 use the formula to find all four solutions of the equation x^4+4x^2+1=0

OpenStudy (jamesj):

Let u = x^2. Then u^2 = x^4 and you can see you now have a quadratic equation in u. Solve that equation for u